# Tony Lindeberg

### v2

@@ 7.1 Variability under uniform scaling transformations
Due to the semi-group property of the affine Gaussian kernel (see Lindeberg (1993) Equation (15.35))
@@
it follows that by combining e.g the output from the receptive fields corresponding to first-order spatial derivatives
@@
where $\nabla_{x}=(\partial_{x_{1}},\partial_{x_{2}})^{T}$, over different positions in image space, we can from the output of such a first layer of visual processing, compute the responses of the first-order derivatives at any coarser level of scale
@@ 7.3 Variability under non-uniform scaling transformations?
In Lindeberg (2024a), an in-depth treatment is given concerning whether the receptive fields in the primary visual cortex could be regarded as spanning a variability over the degree of elongation of the receptive fields.
@@
Unfortunately, there does, however, not appear to exist any sufficiently extensive direct measurements of the eccentricity or the degree of elongation for sufficiently large populations of visual neurons to firmly answer this question. In Lindeberg (2024a), we have therefore taken and alternative indirect approach to this topic, by making use of existing biological measurements of orientation selectivity by Nauhaus et al. (2008), who have reported that there is a substantial variability in the orientation selectivity of the visual receptive fields in relation to the pinwheel structure for monkeys and cats. Goris et al. (2015) have also accumulated histograms of the resultant of the orientation selectivity curves for for simple cells and complex cells.
In Lindeberg (2024b), we have performed an in-depth theoretical analysis of the orientation selectivity properties of idealized models of the receptive fields of simple and complex cells in terms of affine Gaussian derivatives, and shown that there is a direct connection between the orientation selectivity and the degree of elongation of the receptive fields according to the affine Gaussian derivative model. In Lindeberg (2024a) we have furthermore computed closed-form expressions for the resultant of the orientation selectivity curves for the idealized model for visual receptive fields based on affine Gaussian derivatives, and demonstrated that such a closed-form theoretical analysis leads to predictions about histograms of the resultants of orientation selectivity curves for the corresponding idealized models of simple cells that are in very good qualitative agreement with experimentally obtained histograms from biological neurons by Goris et al. (2015).
@@
To more firmly determine whether the hypothesis about an expansion over the degree of elongation of the receptive fields would hold for actual biological neurons, a set of more explicit biological hypotheses, with suggestions for complementary quantitative measurements, have therefore been formulated in Sections 3.2–3.3 in Lindeberg (2024a) to answer this question, and if the working hypothesis would hold, then also characterize how such a possible variability in the degree of elongation of the receptive fields would relate to the pinwheel structures in the orientation maps of the primary visual cortex.
@@ 7.4 Possible variability over a fourth degree of freedom?
Given that a few receptive fields, hence, obviously have been recorded, for which a Gabor-based modelling of receptive fields would be helped by having the principal axes of an affine Gaussian kernel being not oriented in a similar orientation as the principal axis of the cosine function in the Gabor model, one may ask, if when instead modelling those receptive fields by directional derivative of affine Gaussian kernels, such a modelling would be helped by not having the orientation of the directional derivative operator in the affine Gaussian derivative model being aligned with a principal direction of the affine Gaussian kernel used in that affine Gaussian derivative model.
@@ 8 Summary and discussion
In this paper, we have, after describing contextual relations to previous work in Section 2, first in Section 3 performed an in-depth characterization of the degrees of freedom in 2-D spatial affine transformations, based on a decomposition of the affine transformation matrix to a product form very closely related to a singular value decomposition, although on closed form, and with the pre-multiplication and post-multiplication matrices $V$ and $U$ in a regular singular value decomposition here required to be pure rotation matrices.
@@
In these ways, we have described how the idealized affine Gaussian derivative model allows for explicit matching of first of all spatial receptive field responses under general 2-D spatial affine transformations, based on the covariance properties of the spatial affine Gaussian derivative model according to the treatment in Section 4.1. With extension of the spatial affine covariance properties to spatio-temporal affine covariance according to Section 4.2, the proposed model does additionally allow for matching of spatio-temporal receptive field responses computed based on the idealized model for spatio-temporal receptive fields according to (30), when complemented by spatial and temporal differentiation according to the theory presented in Lindeberg (2023; 2024c).
@@
  • If we additionally would interpret the potential support in Lindeberg (2024a) for the hypothesis that the receptive field shapes would additionally be expanded over different degrees of elongations, as predicted from the variabilities of the degree of orientation selectivity established from neurophysiological recordings by Nauhaus et al. (2008) and Goris et al. (2015), it seems natural to also predict that the receptive fields in the primary visual cortex could exhibit properties as would be obtained if the primary visual cortex would have the ability to be covariant also over a significant span over the group of 2-D spatial affine transformations, that is also to image transformations involving non-uniform scaling transformations, as naturally arise from variations of the slant angle, when observing smooth local surface patches in the world from different viewing directions.
To firmly establish if these predictions would hold in reality in the primary visual cortex of higher species, further biological experiments would, however, be extremely valuable, for which explicitly testable experimental hypotheses concerning specific subgroups of 2-D spatial affine transformations have been outlined in Section 3.2.1 in Lindeberg (2023) and Sections 3.2–3.3 in Lindeberg (2024a).

### v3

@@ 7.1 Variability under uniform scaling transformations
Consider the semi-group property of the affine Gaussian kernel (see Lindeberg (1993) Equation (15.35))
@@
from which it follows that by combining e.g the output from the receptive fields corresponding to first-order spatial derivatives
@@
where $\nabla_{x}=(\partial_{x_{1}},\partial_{x_{2}})^{T}$, over different positions in image space. Due to this semi-group property, we can from the output of such a first layer of visual processing compute the responses of the first-order derivatives at any coarser level of scale
@@ 7.3 Variability under non-uniform scaling transformations?
In Lindeberg (2024), an in-depth treatment is given concerning whether the receptive fields in the primary visual cortex could be regarded as spanning a variability over the degree of elongation of the receptive fields.
@@
Unfortunately, there does, however, not appear to exist any sufficiently extensive direct measurements of the eccentricity or the degree of elongation for sufficiently large populations of visual neurons to firmly answer this question. In Lindeberg (2024), we have therefore taken and alternative indirect approach to this topic, by making use of existing biological measurements of orientation selectivity by Nauhaus et al. (2008). They have reported that there is a substantial variability in the orientation selectivity of the visual receptive fields in relation to the pinwheel structure for monkeys and cats. Goris et al. (2015) have also accumulated histograms of the resultant of the orientation selectivity curves for for simple cells and complex cells.
In Lindeberg (2025a), we have performed an in-depth theoretical analysis of the orientation selectivity properties of idealized models of the receptive fields of simple and complex cells in terms of affine Gaussian derivatives, and shown that there is a direct connection between the orientation selectivity and the degree of elongation of the receptive fields according to the affine Gaussian derivative model. In Lindeberg (2024) we have furthermore computed closed-form expressions for the resultant of the orientation selectivity curves for the idealized model for visual receptive fields based on affine Gaussian derivatives, and demonstrated that such a closed-form theoretical analysis leads to predictions about histograms of the resultants of orientation selectivity curves for the corresponding idealized models of simple cells that are in very good qualitative agreement with experimentally obtained histograms from biological neurons by Goris et al. (2015).
@@
To more firmly determine whether the hypothesis about an expansion over the degree of elongation of the receptive fields would hold for actual biological neurons, a set of more explicit biological hypotheses, with suggestions for complementary quantitative measurements, have therefore been formulated in Sections 3.2–3.3 in Lindeberg (2024) to answer this question. If the working hypothesis would hold, the proposed types of neurophysiological measurements could then also characterize how such a possible variability in the degree of elongation of the receptive fields would relate to the pinwheel structures in the orientation maps of the primary visual cortex.
@@ 7.4 Possible variability over a fourth degree of freedom?
Given that a few receptive fields, hence, obviously have been recorded, for which a Gabor-based modelling of receptive fields would be helped by having the principal axes of an affine Gaussian kernel being not oriented in a similar orientation as the principal axis of the cosine function in the Gabor model, one may ask the following: If instead modelling those receptive fields by directional derivative of affine Gaussian kernels, would such modelling be helped by not having the orientation of the directional derivative operator in the affine Gaussian derivative model being aligned with a principal direction of the affine Gaussian kernel used in that affine Gaussian derivative model?
@@
In this context, it is also interesting to note that for some of the receptive field profiles reconstructed by Yazdanbakhsh and Livingstone (2006) (see Figure 6 in that paper) appear to be more similar to first- or second-order directional derivatives of Gaussian kernels in directions different from the principal directions of an affine Gaussian kernel compared to directional derivatives of such kernels in directions that coincide with the principal directions of affine Gaussian kernels. Thus, those reconstructed receptive fields appear to be more similar to members of the receptive field family illustrated in Figure 8 than to the members of the receptive field families illustrated in Figures 77.
@@ 8 Summary and discussion
In this paper, we have given an overall treatment of the variabilities in image structures on the retina or the image plane that are generated by viewing 3-D objects in the environment from different distances and viewing directions, and how the resulting geometric image transformations will interact with the receptive fields in the earliest layers in the visual hierarchy. For convenience of mathematical analysis, we have throughout modelled these geometric image transformations in terms of spatial affine transformations, corresponding to the derivative of the perspective mapping from smooth local surface patches in the world to the retinal surface, or corresponding to the derivative of the projective mappings between pairwise views of the same surface patch from different viewing directions and distances.
For the purpose of establishing an identity relation between receptive field responses computed from different views of the same local surface patch, we have investigated the consequences of the assumption that receptive field family should be covariant under the locally linearized perspective or projective transformations, thereby assuming that the receptive fields should be covariant under spatial affine transformations. As a main working hypothesis, we have thus investigated the consequences of this assumption in the respect of the simple cells in the primary visual cortex being covariant to spatial affine transformations. This covariance property would then mean that the visual receptive fields in the primary visual cortex would meet the variabilities in spatial image structures as generated by the geometric image transformations by a corresponding variability in the receptive field shapes, in order to enable matching of receptive field responses under different viewing conditions of the same object. A manifest implication of that covariance property would then be that the receptive field shapes in the primary visual cortex would span a similar variability as generated by the variability of spatial affine transformations, and thereby span the degrees of freedom of spatial affine transformations.
The main subject of this paper has been to develop a theoretical foundation for investigating the consequences of this way of reasoning, by first relating the degrees of freedom of spatial affine transformations to the degrees of freedom in our highly idealized affine Gaussian derivative model for visual receptive fields, and then considering for which matches of degrees of freedom in the geometric image transformations to the corresponding degrees of freedom in the theoretical receptive field model, we could identify corresponding variabilities in the receptive field shapes of simple cells in the primary visual cortex.
A longer-term underlying motivation to this work is that the presented theoretical foundation could then be used as a basis for more detailed neurophysiological or psychophysical studies, concerning to what extent biological vision for different types species may have developed receptive field structures that have similar properties as the predictions obtained from the highly idealized theory.
More technically, we have, after describing contextual relations to previous work in Section 2, first in Section 3 performed an in-depth characterization of the degrees of freedom in 2-D spatial affine transformations, based on a decomposition of the affine transformation matrix to a product form very closely related to a singular value decomposition, although on closed form, and with the pre-multiplication and post-multiplication matrices $U$ and $V$ in a regular singular value decomposition here required to be pure rotation matrices.
@@
In these ways, we have described how the idealized affine Gaussian derivative model allows for explicit matching of first of all spatial receptive field responses under general 2-D spatial affine transformations, based on the covariance properties of the spatial affine Gaussian derivative model according to the treatment in Section 4.1. With extension of the spatial affine covariance properties to spatio-temporal affine covariance according to Section 4.2, the proposed model does additionally allow for matching of spatio-temporal receptive field responses computed based on the idealized model for spatio-temporal receptive fields according to (30), when complemented by spatial and temporal differentiation according to the theory presented in Lindeberg (2023; 2025b).
@@
  • If we additionally would interpret the potential support in Lindeberg (2024) for the hypothesis that the receptive field shapes would additionally be expanded over different degrees of elongations, as predicted from the variabilities of the degree of orientation selectivity established from neurophysiological recordings by Nauhaus et al. (2008) and Goris et al. (2015). Then, it seems natural to also predict that the receptive fields in the primary visual cortex could exhibit properties as would be obtained if the primary visual cortex would have the ability to be covariant also over a significant span over the group of 2-D spatial affine transformations, that is also to image transformations involving non-uniform scaling transformations, as naturally arise from variations of the slant angle, when observing smooth local surface patches in the world from different viewing directions.
To firmly establish if these predictions, based on our highly idealized theory for visual receptive fields, would hold in reality in the primary visual cortex of higher species, further biological experiments would, however, be extremely valuable, for which explicitly testable experimental hypotheses concerning specific subgroups of 2-D spatial affine transformations have been outlined in Section 3.2.1 in Lindeberg (2023) and Sections 3.2–3.3 in Lindeberg (2024).
@@ 8.1 Outlook
We thus propose the presented theory and predictions to have immediate implications for electrophysiological characterization of receptive fields as well as for psychophysical studies of vision.
One main line of possible future research would concern reconstructing a large set of neurophysiologically recorded receptive field shapes in the primary visual cortex of higher mammals, to map the variabilities in the parameters of the receptive fields (their spatial extent in the principal directions, their orientation, as well as the possible deviation between the direction of a directional derivative operator from the principal directions of the underlying spatial smoothing kernel). In relation to the more detailed proposals concerning such neurophysiological measurements outlined in some additional respects [^3] Note, however, that the treatment in this paper goes conceptually further compared to the earlier treatments in Lindeberg (2023, 2024), by explicitly addressing a variability in receptive field shapes along the fourth degree of freedom in Section 7.4, which was not addressed in Lindeberg (2023, 2024). in Lindeberg (2023, 2024), one may specifically investigate if there would be fundamental differences between lower and higher species, as well as regarding species that either have or not have a pinwheel structure. The developed notion of affine image transformations in relation to receptive field measurements could possibly also be combined with the approach in Yazdanbakhsh and Livingstone (2006) to extract second-order receptive fields, to electrophysiologically characterize the aperture problem, and infer the connectivities in the pathway from the retina, upward to the lateral geniculate nucleus and the primary visual cortex.
Concerning psychophysical experiments, another interesting direction could be to try to explore the influence of evolutionarily- or learning-based priors in the visual system that enable biological vision to very quickly solve visual tasks that from a purely mathematical modelling perspective could be regarded as ambiguous or ill-posed, given that the visual system faces the task of inferring 3-D cues about the environment from lower-dimensional 2-D image data. Could it be established that a pre-wiring of the visual system to achieve constancy of object properties in the environment under variations in the viewing conditions could be characterized in terms of psychophysical measurements, to in turn relate to local affine transformations of image patterns and covariance or invariance properties of the underlying receptive fields under such geometric image transformations, as addressed in this treatment?
For example, regarding visual illusions, the accordion grating illusion, where a non-uniform deformation field perpendicular to the lines is perceived when viewing a set of parallel lines under variations of the distance to the observer, would according to its theoretical explanation (Yazdanbakhsh and Gori 2011) not exist without the existence and influence of local affine image transformations. Additionally, in the rotating tilted lines illusion, where a non-zero radial motion is perceived when viewing a static stimulus consisting of a set of gradually rotated line segments that form a circle at a larger scale, with a geometric interpretation corresponding to the interior of a 3-D tube, can according to its theoretical explanation (Yazdanbakhsh and Gori 2008) be used to estimate the minimum receptive field size as function of the distance from the centre of the visual field.